# Understanding and Mitigating Tokenization Bias in Language Models

### v1

@@ Abstract
State-of-the-art language models are autoregressive and operate on subword units known as tokens. Specifically, one must encode the conditioning string into a list of tokens before passing to the language models for next-token prediction. We show that, for encoding schemes such as maximum prefix matching, tokenization induces a sampling bias that cannot be mitigated with more training or data. To counter this universal problem, we propose a novel algorithm to obtain unbiased estimates from a model that was trained on tokenized data. Our method does not require finetuning the model, and its complexity, defined as the number of model runs, scales linearly with the sequence length. As a consequence, we show that one can simulate token-free behavior from a tokenized language model. We empirically verify the correctness of our method through a Markov-chain setup, where it accurately recovers the transition probabilities, as opposed to the conventional method of directly prompting tokens into the language model.
@@ 1 Introduction
In this work we offer new theoretical insights on the behavior of tokenized LMs. We show that they are statistically equivalent to their token-free counterparts. Specifically, we examine the maximum prefix encoding scheme employed in the WordPiece tokenization method (Devlin et al., 2018; Song et al., 2020) and find that this process not only results in biased estimates of next token probabilities, but also leads to overall skewed estimates of subsequent character probabilities. In general, this bias persists despite an increase in training data, even within the simple setting of a 1st-order Markov chain. Such bias occurs due to the implicit disparity between the domain of the conditioning context, namely, characters versus tokens. Nevertheless, we will show that it is possible to correct this bias without resorting to finetuning. Once adjusted, it becomes possible to simulate the token-free behavior learned implicitly by the tokenized LM and even (theoretically) mimic the behavior of another tokenized model employing a distinct vocabulary set, all without requiring finetuning. Our specific contributions are as follows:
@@
  • We present a novel algorithm capable of transforming a tokenized LM into its token-free counterpart, thereby effectively eliminating the sampling bias.
  • We verify the correctness of our algorithm on learning the transition matrix of a $k$-th order Markov chain.
@@ 3 $AA$
Tokenization Setting. We assume having a predefined vocabulary $\mathcal{V}$ constructed using any tokenization algorithm such as Byte-Pair-Encoding (BPE) or Lempel-Ziv-Welch (LZW), with the condition that $\mathcal{A}\subseteq\mathcal{V}$. We use $t$ to denote a token in $\mathcal{V}$, i.e. $t\in\mathcal{V}$. Importantly, we use the longest prefix matching strategy for tokenization (encoding), denoted as $\mathrm{encode}(.)$, similar to the approach used in the Wordpiece algorithm (Devlin et al., 2018; Song et al., 2020). Given a sequence of tokens $t^{k}_{1}$, the function $\mathrm{decode}(.)$ returns the concatenated string resulting from processing each token in the sequence. Finally, the set of all strings that starts with the tokens $t^{k}_{1}$ is defined as $\mathcal{S}(t^{k}_{1}){=}\{s|t^{k}_{1}{=}\mathrm{encode}(s)^{k}_{1}\}$.
@@ 2.2 Next-Character Sampling Bias
Since this applies to any optimally trained LM, increasing the training set size does not mitigate this problem. The reason for this sampling bias is that, during the tokenization process with longest prefix matching, the token ``$A"$ must be followed by the token ``$B"$. Otherwise, the encoding process will merge to create a longer token ``$AA"$. To generalize this phenomenon, we start with the definition of invalid encodings.
@@
(Token-Induced Zero Probability) Let $t^{i}_{1}$ be a sequence of input tokens. For any invalid encoding $t^{i}_{1}$, we have $P_{\mathrm{gt}}(t^{i}_{1}){=}0.0$ and the conditional probability $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1})$ is undefined. In the case $t^{i}_{1}$ is valid, then $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1}){=}0.0$ if $t^{i+1}_{1}$ is invalid. Furthermore, let $x^{n}_{1}{=}\mathrm{decode}(t^{i}_{1})$, then for any string $x^{N}_{n+1}$ such that $\mathrm{encode}(\mathrm{concat}(\mathrm{decode}(t^{i}_{1}),x^{N}_{n+1})){\neq}t^{i}_{1}$, then $P_{\mathrm{gt}}(x^{N}_{n+1}|t^{i}_{1}){=}0.0.$
@@ 3 Alleviating Sampling Bias
Our method consists of two stages. In the first stage, the idea is to identify the condition when $P(x^{N}_{n+1}|t^{i}_{1})=P(x^{N}_{n+1}|x^{n}_{1})$ where $t^{i}_{1}=\mathrm{encode}(x^{n}_{1})$. Once identified, we can refactor the conditional probability to match the conditioning events. In the second stage, we compute $P(x^{N}_{n+1}|t^{i}_{1})$ using the LM output probability, i.e. $P(t_{i+1}|t^{i}_{1})$, through the novel Branch and Pass (BPTree) Algorithm.
@@ 3.1 Refactoring
Our method removes the bias by connecting character and token domains through a special subset of tokens $\mathcal{V}^{*}{\subset}\mathcal{V}$, whose elements $t^{*}{\in}\mathcal{V}^{*}$ are not a substring of any other tokens in $\mathcal{V}$ but itself. For example, given $\mathcal{V}{=}\{``AAA",``AA",``CB",``A",``B",``C"\}$, then $\mathcal{V}^{*}=\{``AAA",``CB"\}$. In the Markov example in Section 2, this corresponds to the tokens ``$AA"$ and ``$B"$. Also, we assume that any string $x^{N}_{1}$ has the first token $t_{1}\in\mathcal{V}^{*}$ [^3] This is not restrictive, as many current language models starts with a special beginning token in $\mathcal{V}^{*}$, e.g. the start token ${<}start{>}$ in SentencePiece (Kudo & Richardson, 2018).. Consider the input string $x^{n}_{1}$ and its corresponding encoding $t^{i}_{1}{=}\mathrm{encode}(x^{n}_{1})$, Proposition 3.1 shows the sufficient condition for $\mathcal{S}(t^{i}_{1}){=}\mathcal{S}(x^{n}_{1})$.
@@
where $k$ is the last token in $\mathrm{encode}(x^{n}_{1})$ such that $t_{k}{\in}\mathcal{V}^{*}$ and $x^{n_{k}}_{1}{=}\mathrm{decode}(t^{k}_{1})$, where $n_{k}\leq n$. Proof details of this step can be found in the Appendix E. We then use the BPTree algorithm to compute each term in the RHS individually.
3.2 The Branch and Pass Algorithm
Algorithm 1 Branch and Pass Algorithm on Discrete Probability Tree. This algorithm recursively computes $P(x^{N}_{n_{k}+1}|t^{k}_{1})$.
1:procedure BPTree($x^{N}_{n_{k}+1},t^{k}_{1}$)
@@ 3.2 The Branch and Pass Algorithm
13:    $\mathrm{p_{val}}=\mathrm{p_{val}}\times\text{BPTree}(x^{N}_{n_{k+1}+1},t^{k+1}_{1})$
@@
We present the BPTree algorithm in Algorithm 1, that allows us to compute the probabilities $P(x^{N}_{n_{k}+1}|t^{k}_{1})$ and $P(x^{n}_{n_{k}+1}|t^{k}_{1})$ in Equation (1). Note that this algorithm does not require $t_{k}{\in}\mathcal{V}^{*}$. Details on the algorithmic correctness are shown in Appendix E.
The idea is to marginalize out $P(x^{N}_{n_{k}+1}|t^{k}_{1})$ by considering two complementary events: when the next token $t_{k+1}$ has a prefix $x^{N}_{n_{k}+1}$ ($\mathrm{b_{val}}$ in the Branch Step) versus when the next token $t_{k+1}$ is contained within $x^{N}_{n_{k}+1}$ ($\mathrm{p_{val}}$ in the Pass Step). Formally, BPTree computes the following probabilities:
@@
We provide an intuitive explanation for the algorithm following the example in Figure 2. Here, we would like to compute the probability $P(x^{n_{k}+3}_{n_{k}+1}{=}``bee"|t^{k}_{1})$. The first possibility is that ``$bee"$ is a prefix of the next token, so we search for all such tokens (line 3 in the algorithm) and sum up their probability (line 4), i.e. $\mathrm{b_{val}}{=}P(t_{k+1}{=}``beer"|t^{k}_{1})$. Figure 2 visualizes this step as branching out the tree by finding all tokens completing the string. Since ``$beer"$ is not the only string that contains ``$bee"$, e.g. ``$beep",``been"$, etc. we need to compute the probability for these other scenarios, each of which has $t_{k+1}{=}``b"$ (the first token in ``$bee"$, line 10 and 12) due to maximum prefix encoding. Then, we want to compute the probability that the subsequent string is ``$ee"$ (line 13), given the previous $t^{k}_{1}$ and $t_{k+1}{=}``b"$, which is the output of the BPTree algorithm but for $x^{n_{k}+3}_{n_{k}+2}{=}``ee"$ and $t^{k+1}_{1}$. Formally, in the Passing step: $\mathrm{p_{val}}{=}P(t_{k+1}{=}``b"|t^{k}_{1})P(x^{n_{k}+3}_{n_{k}+2}{=}``ee"|t^{k}_{1},t_{k+1}{=}``b")$. We continue the procedure until meeting the base case, where the string must be a prefix of the next token (usually, when there is only a single character left). Finally, by computing the sum of the branch and pass steps, we obtain the desired conditional probability $\mathrm{b_{val}}{+}\mathrm{p_{val}}{=}P(x^{n_{k}+3}_{n_{k}+1}{=}``bee"|t^{k}_{1})$.
@@ 4 Experiments
Figure 2: BPTree Visualization. At each recursive call, the Branch step finds tokens that starts with the query string while the Pass step extracts and employs the next token and leftover string for the next recursive call until meeting the base case.
@@ 5 Conclusion
This work identifies the next-character sampling gap between a tokenized model and a token-free one, which persists even for optimally trained models. We present a probabilistic approach to effectively eliminate this bias without requiring additional training. This closes the sampling gap between tokenized and token-free models, suggesting that language models implicitly absorb character-level information despite being trained solely on tokenized text. This result implies that it is theoretically possible to simulate the behavior of another language model trained using different vocabulary without any fine-tuning, since it is possible to transfer from token-free models to tokenized counterparts. Future work seeks to extend our investigation for alternative encoding strategies, most notably the widely used BPE tokenization (Sennrich et al., 2015), in state-of-the-art models.
@@ Appendix A Related Work
Appendix B Supporting Theorems on Tokenization
This section provides supporting theorems for the proof of the main results. We first remind the readers that the set $\mathcal{S}(x^{n}_{1})$ corresponds to the set of all strings that contains $x^{n}_{1}$ as a prefix. Similarly, the event set $\mathcal{S}(t^{i}_{1})$ corresponds to the set of all strings whose first $i$ tokens are $t^{i}_{1}$. Consider when $t^{i}_{1}=\mathrm{encode}(x^{n}_{1})$, it should be noted that the two sets $S(t^{i}_{1})$ and $S(x^{n}_{1})$ are not guaranteed to be equivalent. That is because the subsequent characters after $x^{n}_{1}$ can affect the tokenization within the first $n$ character. We illustrate this in more details in the following example.
@@ Appendix B Supporting Theorems on Tokenization
We introduce the following proposition that contains two facts regarding the maximum prefix encoding process.
@@
  • If $t^{k}_{1}$ is invalid then $\mathcal{S}(t^{k}_{1})=\emptyset$: We prove this by showing that it is impossible to have a string with prefix $\mathrm{decode}(t^{k}_{1})$ whose first $k$ tokens are $t^{k}_{1}$. Let $\bar{s}=\mathrm{decode}(t^{k}_{1})$, we have any string $s\in\mathcal{S}(t^{k}_{1})$ must have $\bar{s}\in\mathrm{prefix}(s)$. Assume such string $s$ exists, we have $\mathrm{encode}(s)^{k}_{1}=t^{k}_{1}$. Let $x^{n}_{1}=\mathrm{decode}(t^{k}_{1})$, according to Proposition B.1, we have $r=\bar{s}$ and as the result, for $1\leq i\leq k$, we have $\mathrm{encode}(s)_{i}=t_{i}=\mathrm{encode}(x^{n}_{1})_{i}$, which contradicts our initial assumption. This completes the proof.
@@
Proposition 2.3 (Token-Induced Zero Probability) Let $t^{i}_{1}$ be a sequence of input tokens. For any invalid encoding $t^{i}_{1}$, we have $P_{\mathrm{gt}}(t^{i}_{1}){=}0.0$ and the conditional probability $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1})$ is undefined. In the case $t^{i}_{1}$ is valid, then $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1}){=}0.0$ if $t^{i+1}_{1}$ is invalid. Furthermore, let $x^{n}_{1}{=}\mathrm{decode}(t^{i}_{1})$, then for any string $x^{N}_{n+1}$ such that $\mathrm{encode}(\mathrm{concat}(\mathrm{decode}(t^{i}_{1}),x^{N}_{n+1})){\neq}t^{i}_{1}$, then $P_{\mathrm{gt}}(x^{N}_{n+1}|t^{i}_{1}){=}0.0.$
@@
    Let $s^{\prime}=\mathrm{decode}(t^{i}_{1})$, note that $P_{\mathrm{gt}}(x^{N}_{n+1},t^{i}_{1})=P_{\mathrm{gt}}(s^{\prime},x^{N}_{n+1},t^{i}_{1})=P_{\mathrm{gt}}(\bar{s},t^{i}_{1})$ where $\bar{s}=\mathrm{concat}(s^{\prime},x^{N}_{n+1})=x^{N}_{1}$.
@@
    Consider $P_{\mathrm{gt}}(\bar{s},t^{i}_{1})=P_{\mathrm{gt}}(\bar{s})P_{\mathrm{gt}}(t^{i}_{1}|\bar{s})$, we will prove that $P_{\mathrm{gt}}(t^{i}_{1}|\bar{s})=0.0$ if $\mathrm{encode}(\bar{s})^{i}_{1}\neq t^{i}_{1}$.
The idea is that, when we append additional characters to $\bar{s}$, it will never change the tokenization such that the first $i$ tokens are $t^{i}_{1}$. Formally:
  • Let $j$ be the first position such that $\mathrm{encode}(\bar{s})_{j}\neq t_{j}$ then we know that $|\mathrm{decode}(\mathrm{encode}(\bar{s})_{j}))|>|\mathrm{decode}(t_{j})|$ (Proposition B.1 (Result 2)).
  • Following Proposition B.1 (Result 2), let $s^{*}\in\mathcal{S}(\bar{s})$, then we know that $\mathrm{decode}(\mathrm{encode}(\bar{s})_{j})$ must be a substring of within another longer token (it cannot be broken down). This completes the proof.
@@
1) General Case: Since there exists a string $s$ with prefix $x^{n}_{1}$ where $\mathrm{encode}(s)^{i}_{1}\neq t^{i}_{1}$, which follows directly from our 1st order Markov chain example in the main paper, i.e. the string $s=``AA"$ has ``$A"$ as prefix but have the $t_{1}=``AA"\neq``A"$. On the other hand, we know that any string $s$ that has the first $i$ tokens as $t^{i}_{1}$ must have the first $n$ characters as $x^{n}_{1}$, as such $\mathcal{S}(t^{i}_{1})\subset\mathcal{S}(x^{n}_{1})$ and $P(x^{n}_{1}|t^{i}_{1})=1.0$.
@@
Remarks. We briefly note that the condition $t_{i}\in\mathcal{V}^{*}$ is the sufficient condition. In general, any token sequence $t^{i}_{1}$ that satisfies $\mathcal{S}(t^{i}_{1})=\mathcal{S}(x^{n}_{1})$ will have $P(t^{i}_{1}|x^{n}_{1})=1.0$. Developing an efficient algorithm to verify whether $\mathcal{S}(t^{i}_{1})=\mathcal{S}(x^{n}_{1})$ remains a task for future research.

### v2

@@ Abstract
State-of-the-art language models are autoregressive and operate on subword units known as tokens. Specifically, one must encode the conditioning string into a list of tokens before passing to the language models for next-token prediction. We show that popular encoding schemes, such as maximum prefix encoding (MPE) and byte-pair-encoding (BPE), induce a sampling bias that cannot be mitigated with more training or data. To counter this universal problem, for each encoding scheme above, we propose a novel algorithm to obtain unbiased estimates from any language model trained on tokenized data. Our methods do not require finetuning the model, and the complexity, defined as the number of model runs, scales linearly with the sequence length in the case of MPE. As a result, we show that one can simulate token-free behavior from a tokenized language model. We empirically verify the correctness of our method through a Markov-chain setup, where it accurately recovers the transition probabilities, as opposed to the conventional method of directly prompting tokens into the language model.
@@ 1 Introduction
In this work we offer new theoretical insights on the behavior of tokenized LMs. We show that they are statistically equivalent to their token-free counterparts. Specifically, we examine the maximum prefix encoding (MPE) scheme employed in the WordPiece tokenization method (Devlin et al., 2018; Song et al., 2020) and find that this process not only results in biased estimates of next token probabilities, but also leads to overall skewed estimates of subsequent character probabilities. In general, this bias persists despite an increase in training data, even within the simple setting of a 1st-order Markov chain. Such bias occurs due to the implicit disparity between the domain of the conditioning context, namely, characters versus tokens. Nevertheless, we will show that it is possible to correct this bias without resorting to finetuning. Once adjusted, it becomes possible to simulate the token-free behavior learned implicitly by the tokenized LM and even (theoretically) mimic the behavior of another tokenized model employing a distinct vocabulary set, all without requiring finetuning. Our specific contributions are as follows:
@@
  • We present two novel algorithms to correct this bias for MPE and Byte-Pair-Encoding (BPE) respectively. Due to space limit, the analysis and algorithm for BPE are presented in Appendix H.
  • We verify the correctness of our algorithms on learning the transition matrix of a $k$-th order Markov chain.
@@ 3 $AA$
Tokenization Setting. We assume having a predefined vocabulary $\mathcal{V}$ constructed using any tokenization algorithm such as BPE, with the condition that $\mathcal{A}{\subseteq}\mathcal{V}$. We use $t$ to denote a token in $\mathcal{V}$, i.e. $t{\in}\mathcal{V}$. Importantly, we use the longest prefix matching strategy for tokenization (encoding), denoted as $\mathrm{encode}(.)$, similar to the approach used in the Wordpiece algorithm (Devlin et al., 2018; Song et al., 2020). Given a sequence of tokens $t^{k}_{1}$, the function $\mathrm{decode}(.)$ returns the concatenated string resulting from processing each token in the sequence. Finally, the set of all strings that starts with the tokens $t^{k}_{1}$ is defined as $\mathcal{S}(t^{k}_{1}){=}\{s|t^{k}_{1}{=}\mathrm{encode}(s)^{k}_{1}\}$.
@@ 2.2 Next-Character Sampling Bias
Since this applies to any optimally trained LM, increasing the training set size does not mitigate this problem. The reason for this sampling bias is that, during the tokenization process with longest prefix matching, the token ``$A"$ must be followed by the token ``$B"$. Else, MPE encoding will merge to create a longer token ``$AA"$. We generalize this phenomenon with the definition of invalid encodings.
@@
(Token-Induced Zero Probability) Let $t^{i}_{1}$ be a sequence of input tokens. For any invalid encoding $t^{i}_{1}$, we have $P_{\mathrm{gt}}(t^{i}_{1}){=}0.0$ and the conditional probability $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1})$ is undefined. In the case $t^{i}_{1}$ is valid, then $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1}){=}0.0$ if $t^{i+1}_{1}$ is invalid. Furthermore, let $x^{n}_{1}{=}\mathrm{decode}(t^{i}_{1})$, then for any string $x^{N}_{n+1}$ such that $\mathrm{encode}(\mathrm{concat}(\mathrm{decode}(t^{i}_{1}),x^{N}_{n+1})){\neq}t^{i}_{1}$, we have $P_{\mathrm{gt}}(x^{N}_{n+1}|t^{i}_{1}){=}0.0.$
@@ 3 Alleviating Sampling Bias
Our method consists of two stages. In the first stage, the idea is to identify the condition when $P(x^{N}_{n+1}|t^{i}_{1})=P(x^{N}_{n+1}|x^{n}_{1})$ where $t^{i}_{1}=\mathrm{encode}(x^{n}_{1})$. Once identified, we can refactor the conditional probability to match the conditioning events. In the second stage, we compute $P(x^{N}_{n+1}|t^{i}_{1})$ using the LM output probability, i.e. $P(t_{i+1}|t^{i}_{1})$, through the novel Maximum Prefix Correction (MPC) Algorithm.
@@ 3.1 Refactoring
Our method removes the bias by connecting character and token domains through a special subset of tokens $\mathcal{V}^{*}{\subset}\mathcal{V}$, whose elements $t^{*}{\in}\mathcal{V}^{*}$ are not a substring of any other tokens in $\mathcal{V}$ but itself. For example, given $\mathcal{V}{=}\{``AAA",``AA",``CB",``A",``B",``C"\}$, then $\mathcal{V}^{*}=\{``AAA",``CB"\}$. In the Markov example in Section 2, this corresponds to the tokens ``$AA"$ and ``$B"$. Also, we assume that any string $x^{N}_{1}$ has the first token $t_{1}\in\mathcal{V}^{*}$ [^3] Many current language models begins with a start token ${<}\mathrm{start}{>}$ in $\mathcal{V}^{*}$, e.g. in SentencePiece (Kudo & Richardson, 2018).. Consider the input string $x^{n}_{1}$ and its corresponding encoding $t^{i}_{1}{=}\mathrm{encode}(x^{n}_{1})$, Proposition 3.1 shows the sufficient condition for $\mathcal{S}(t^{i}_{1}){=}\mathcal{S}(x^{n}_{1})$.
@@
where $k$ is the last token in $\mathrm{encode}(x^{n}_{1})$ such that $t_{k}{\in}\mathcal{V}^{*}$ and $x^{n_{k}}_{1}{=}\mathrm{decode}(t^{k}_{1})$, where $n_{k}\leq n$. Proof details of this step can be found in the Appendix E. We then use the MPC algorithm to compute each term in the RHS individually.
3.2 Maximum Prefix Correction Algorithm
Algorithm 1 Maximum Prefix Correction Algorithm. This algorithm recursively computes $P(x^{N}_{n_{k}+1}|t^{k}_{1})$.
1:procedure compute($x^{N}_{n_{k}+1},t^{k}_{1}$)
@@ 3.2 Maximum Prefix Correction Algorithm
13:    $\mathrm{p_{val}}=\mathrm{p_{val}}\times\text{COMPUTE}(x^{N}_{n_{k+1}+1},t^{k+1}_{1})$
@@
We present the MPC algorithm in Algorithm 1, that allows us to compute the probabilities $P(x^{N}_{n_{k}+1}|t^{k}_{1})$ and $P(x^{n}_{n_{k}+1}|t^{k}_{1})$ in Equation (1). Note that this algorithm does not require $t_{k}{\in}\mathcal{V}^{*}$. Details on the algorithmic correctness are shown in Appendix E.
The idea is to marginalize out $P(x^{N}_{n_{k}+1}|t^{k}_{1})$ by considering two complementary events: when the next token $t_{k+1}$ has a prefix $x^{N}_{n_{k}+1}$ ($\mathrm{b_{val}}$ in the Branch Step) versus when the next token $t_{k+1}$ is contained within $x^{N}_{n_{k}+1}$ ($\mathrm{p_{val}}$ in the Pass Step). Formally, MPC computes the following probabilities:
@@
We provide an intuitive explanation for the algorithm following the example in Figure 2. Here, we would like to compute the probability $P(x^{n_{k}+3}_{n_{k}+1}{=}``bee"|t^{k}_{1})$. The first possibility is that ``$bee"$ is a prefix of the next token, so we search for all such tokens (line 3 in the algorithm) and sum up their probability (line 4), i.e. $\mathrm{b_{val}}{=}P(t_{k+1}{=}``beer"|t^{k}_{1})$. Figure 2 visualizes this step as branching out the tree by finding all tokens completing the string. Since ``$beer"$ is not the only string that contains ``$bee"$, e.g. ``$beep",``been"$, etc. we need to compute the probability for these other scenarios, each of which has $t_{k+1}{=}``b"$ (the first token in ``$bee"$, line 10 and 12) due to maximum prefix encoding. Then, we want to compute the probability that the subsequent string is ``$ee"$ (line 13), given the previous $t^{k}_{1}$ and $t_{k+1}{=}``b"$, which is the output of the MPC algorithm but for $x^{n_{k}+3}_{n_{k}+2}{=}``ee"$ and $t^{k+1}_{1}$. Formally, in the Passing step: $\mathrm{p_{val}}{=}P(t_{k+1}{=}``b"|t^{k}_{1})P(x^{n_{k}+3}_{n_{k}+2}{=}``ee"|t^{k}_{1},t_{k+1}{=}``b")$. We continue the procedure until meeting the base case, where the string must be a prefix of the next token (usually, when there is only a single character left). Finally, by computing the sum of the branch and pass steps, we obtain the desired conditional probability $\mathrm{b_{val}}{+}\mathrm{p_{val}}{=}P(x^{n_{k}+3}_{n_{k}+1}{=}``bee"|t^{k}_{1})$.
@@ 4 Experiments
Figure 2: MPC Visualization. At each recursive call, the Branch step finds tokens that starts with the query string while the Pass step extracts and employs the next token and leftover string for the next recursive call until meeting the base case.
@@ 5 Conclusion
This work identifies the next-character sampling gap between a tokenized model and a token-free one, which persists even for optimally trained models. We present a probabilistic approach to effectively eliminate this bias without requiring additional training. This closes the sampling gap between tokenized and token-free models, suggesting that language models implicitly absorb character-level information despite being trained solely on tokenized text. This result implies that it is theoretically possible to simulate the behavior of another language model trained using different vocabulary without any fine-tuning, since it is possible to transfer from token-free models to tokenized counterparts.
@@ Appendix A Related Work
Appendix B Supporting Theorems on Maximum Prefix Encoding
This section provides supporting theorems for the proof of the main results. We first remind the readers that the set $\mathcal{S}(x^{n}_{1})$ corresponds to the set of all strings that contain $x^{n}_{1}$ as a prefix. Similarly, the event set $\mathcal{S}(t^{i}_{1})$ corresponds to the set of all strings whose first $i$ tokens are $t^{i}_{1}$. Consider when $t^{i}_{1}=\mathrm{encode}(x^{n}_{1})$, it should be noted that the two sets $S(t^{i}_{1})$ and $S(x^{n}_{1})$ are not guaranteed to be equivalent. That is because the subsequent characters after $x^{n}_{1}$ can affect the tokenization within the first $n$ character. We illustrate this in more detail in the following example.
@@ Appendix B Supporting Theorems on Maximum Prefix Encoding
Figure 4: Interpretations of Proposition B.1, which shows that for any string $s$ with prefix $x^{n}_{1}$, its token within $x^{n}_{1}$ must start at a certain designated positions. For each encoding, same color denotes belonging to the same token. This would later allows us to construct an efficient algorithm to correct the bias. See Corollary B.3 for details on invalid encoding.
We introduce the Proposition B.1 that contains two facts regarding the MPE process, visually presented in Figure 4.
@@
  • If $t^{k}_{1}$ is invalid then $\mathcal{S}(t^{k}_{1})=\emptyset$: Let $x^{n}_{1}=\mathrm{decode}(t^{k}_{1})$ and $t^{i}_{1}=\mathrm{encode}(x^{n}_{1})$. Let $s_{1}\in\mathcal{S}(t^{i}_{1})$ and suppose there exist a string $s_{2}\in\mathcal{S}(t^{k}_{1})$. Re-running the MPE procedure on $s_{1}$ and $s_{2}$ in parallel, then every time a token is selected within $x^{n}_{1}$ in $s_{1}$, it must also be selected at the same position in $s_{2}$ as well. Thus, we cannot have $t^{i}_{1}\neq t^{k}_{1}$, which proves the result.
@@
Proposition 2.3 (Token-Induced Zero Probability) Let $t^{i}_{1}$ be a sequence of input tokens. For any invalid encoding $t^{i}_{1}$, we have $P_{\mathrm{gt}}(t^{i}_{1}){=}0.0$ and the conditional probability $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1})$ is undefined. In the case $t^{i}_{1}$ is valid, then $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1}){=}0.0$ if $t^{i+1}_{1}$ is invalid. Furthermore, let $x^{n}_{1}{=}\mathrm{decode}(t^{i}_{1})$, then for any string $x^{N}_{n+1}$ such that $\mathrm{encode}(\mathrm{concat}(\mathrm{decode}(t^{i}_{1}),x^{N}_{n+1})){\neq}t^{i}_{1}$, we have $P_{\mathrm{gt}}(x^{N}_{n+1}|t^{i}_{1}){=}0.0.$
@@
    Note that $P_{\mathrm{gt}}(x^{N}_{n+1},t^{i}_{1})=P_{\mathrm{gt}}(x^{N}_{1},t^{i}_{1})$ where $\mathrm{concat}(x^{n}_{1},x^{N}_{n+1})=x^{N}_{1}$.
@@
    Consider $P_{\mathrm{gt}}(x^{N}_{1},t^{i}_{1})=P_{\mathrm{gt}}(x^{N}_{1})P_{\mathrm{gt}}(t^{i}_{1}|x^{N}_{1})$, we will prove that $P_{\mathrm{gt}}(t^{i}_{1}|x^{N}_{1})=0.0$ if $\mathrm{encode}(x^{N}_{1})^{i}_{1}\neq t^{i}_{1}$.
The proof idea for this is shown in Figure 4 (Example 2, Right). Formally:
  • Let $j$ be the first position such that $\mathrm{encode}(x^{N}_{1})_{j}\neq t_{j}$ then we know that $|\mathrm{decode}(\mathrm{encode}(x^{N}_{1})_{j}))|>|\mathrm{decode}(t_{j})|$ (Proposition B.1 (Result 2)).
  • Following Proposition B.1 (Result 2), let $s\in\mathcal{S}(x^{N}_{1})$, then we know that $\mathrm{decode}(\mathrm{encode}(x^{N}_{1})_{j})$ must be a substring of within another longer token (it cannot be broken down) in $s$. Hence, no string $s$ will have a $j$-th token as $t_{j}$, so $P_{\mathrm{gt}}(t^{i}_{1}|x^{N}_{1})=0.0$. This completes the proof.
@@
1) General Case: There exists a string $s\in\mathcal{S}(x^{n}_{1})$ where $\mathrm{encode}(s)^{i}_{1}\neq t^{i}_{1}$, following directly from our 1st order Markov chain example in the main paper, i.e. the string $s=``AA"$ has ``$A"$ as prefix but have the $t_{1}=``AA"\neq``A"$. Also, any string $s$ that has the first $i$ tokens as $t^{i}_{1}$ must have the first $n$ characters as $x^{n}_{1}$, hence $\mathcal{S}(t^{i}_{1})\subset\mathcal{S}(x^{n}_{1})$ and $P(x^{n}_{1}|t^{i}_{1})=1.0$.
@@
Remarks. We briefly note that the condition $t_{i}\in\mathcal{V}^{*}$ is the sufficient condition. In general, any token sequence $t^{i}_{1}$ that satisfies $\mathcal{S}(t^{i}_{1})=\mathcal{S}(x^{n}_{1})$ will have $P(t^{i}_{1}|x^{n}_{1})=1.0$. One potential strategy is to find the first index $i=0,1,...k-1$ such that $t^{k}_{k-i}$ cannot be merged into another token in $\mathcal{V}$.